• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name orchen

    Log entry time 17:57:55 on May 08, 2011

    Entry number 358642

    keyword=HAND recovery problems

    We have two problems. (1) Left PMs of bars 1,2,4,10,11 from layer 6 are dead. We will try bringing them back tomorrow and see if they are stabelising. We traced down the problem to the PMs them self so we can't do much about it now.

    (2) We have events in the ADC spectra that are below the padestal. This is common when runing high current beam but we never got this with cosmic before.

    We have no clear conclutions on this issue. We will have to see how it looks like with beam.

    The figures below are:

    1. TDC spectra. As can be seen we have the expected peak at ** but we also get a second peak at **. This is common to all PMs. It is probbly related to problems in the PMs arissing from the high current runing of x>2. Nothig we tried to do inorder to fix it helped.

    2. ADC specta. Black: No cuts Red: with a cut on TDC>0 Green: With a cut on 0<TDC<1500

    The events that are below the padestal are coming from the other peak in the TDC spectra.



    A copy of this log entry has been emailed to: eip@tauphy.tau.ac.il , sgilad@mit.edu , doug@jlab.org , vasulk@jlab.org , igor.korover@gmail.com , rshneor@jlab.org , israel_yaron@yahoo.com , sbeck@bgu.ac.il



    Figure 1



    Figure 2